I am unsure of what to do with this problem because there is a negative sign in front of some of the variables.
-x^6*x^2*-x^12

Answer by sdmmadam@yahoo.com(530)

If it is minus x we are talking about and then the poweris 6,or 12 or what ever then the problem ought to have been given as
(-x)^6 X(x^2)X (-x)^12
=(x)^6 X(x^2)X(x)^12
=(x)^(6+2+12) (while quantities of the same base are multiplied,write the base and add the powers is the rule)
=x^20
( why? what has happened?have we simply left out the not so pleasant minus sign without rhyme or reason? No.Not at all. We have used the mathematical concept that (-1) raised to any even integer power is (+1)
Here (-x) means (-1)X(x) and
(-x)^6 means [(-1)X(x)]^6 (using (ab)^m = (a^m)(b^m) )
= (-1)^6X(x)^6 = 1X(x^6) = x^6
Supposing the problem is
-(x^6)X(x^2)X[-(x^12)] then since (negativeXpositiveXnegative = positive)
=+(x)^(6+2+12)
=x^20